# Coupled dimerized alternating-bond quantum spin chains in the distorted honeycomb-lattice magnet Cu5SbO6

### v1

@@ Abstract
We analyze powder-averaged inelastic neutron scattering and magnetization data for the distorted honeycomb compound Cu5SbO6 using a first-order dimer expansion calculation and quantum Monte Carlo simulations. We show that, in contrast to the previously proposed honeycomb lattice model, Cu5SbO6 accommodates interacting dimerized spin chains with alternating ferromagnetic-antiferromagnetic couplings along the chain. Moreover, unlike the typical couplings observed in other Cu2+-based distorted honeycomb magnets, the spin chains in Cu5SbO6 primarily couple through an antiferromagnetic coupling $J_{4}$ that arises between the honeycomb layers, rather than the expected interchain $J_{3}$ coupling in the layers. This finding reveals a different magnetic coupling scheme, $J_{1}$-$J_{2}$-$J_{4}$, for Cu5SbO6. In addition, utilizing x-ray spectroscopy and transmission electron microscopy, we also refine the crystal structure and stacking-fault model of the compound.
@@ I Introduction
Dimerized quantum magnets have constituted a fertile ground for exploring a plethora of exotic emergent quantum phenomena in magnetic insulators [Sachdev2008, Vasiliev2018]. The most renowned aspect of such magnets is arguably quantum criticality. One of the quantum phase transitions is achieved in the field-induced Bose-Einstein condensation of magnons realized in several compounds [Nikuni1999, Ruegg2003, Jaime2004, Zapf2006, Giamarchi2008, Aczel2009, Zapf2014, Samulon2009]. Novel phenomena in the vicinity of a quantum critical point separating a dimerized disordered phase from a long-range ordered phase such as dimensional reduction [Sebastian2006] and longitudinal excitation modes [Ruegg2008] have been observed. When frustration between exchange couplings among orthogonally oriented dimers are present as in the Shastry-Sutherland magnet, several other intriguing effects have been reported including topologically protected chiral edge modes of triplet excitations [McClarty2017], pressure-induced 4-spin plaquette singlet state [Koga2000, Zayed2017], field-induced spin-nematic phase [Fogh2024], incomplete devil’s staircase [Takigawa2013] and crystallization of triplet bound states in magnetization plateaus [Corboz2014]. Dimerized magnets are built upon spin dimers consisting of a pair of spins interacting with an intra-dimer antiferromagnetic (AFM) exchange coupling. The dimer, for the case of $S=\frac{1}{2}$, exhibits a non-magnetic singlet ground state and degenerate triplet excited states in zero magnetic field which are separated by a finite energy gap. Dimers that interact with each other with a ferromagnetic (FM) exchange coupling in a one-dimensional (1D) network, forming an alternating FM-AFM spin chain, form another distinct class of dimerized magnets. Magnetic materials that are characterized as quasi-1D FM-AFM dimerized spin chains include, as examples, copper nitrate [Tennant2012], BaCu2V2O8 [Klyushina2018], CuNb2O6 [Kodama1999], (CH3)2NH2CuCl3 [Stone2007], etc.
@@
Figure 1: (a) A monoclinic unit cell of Cu5SbO6 is displayed in which spin chains consisted of SbO6 octahedra (in silver) and edge-sharing Cu2O6 double plaquettes (in green) are aligned along the $a$-axis. The unit cell boundary is indicated by the dashed lines. (b) A chain segment with alternating $J_{1}$ and $J_{2}$ couplings is shown. The dashed purple and solid yellow lines represent intra-dimer $J_{1}$ and inter-dimer $J_{2}$ couplings, respectively. (c) The abab-plane projection of a single Cu2SbO6 layer is displayed. $B_{1}$ = 2.314 Å is the Cu-O bond perpendicular to the plaquette whereas $B_{2}$ = 1.994 Å and $B_{3}$ = 2.043 Å are the Cu-O bonds within the plaquette. The dot-dash turquoise lines represent inter-chain $J_{3}$ couplings. (d) A single layer projected on to the acac-plane is shown, where the dotted orange lines represent inter-chain $J_{4}$ couplings perpendicular to the honeycomb planes. The super-exchange path responsible for the $J_{4}$ coupling in Cu5SbO6 is displayed in (e) along with the corresponding path in Na2Cu2TeO6. The crystal structure illustrations are generated by VESTA [Momma2011].
@@
To our knowledge, the archetypes of the quasi-1D $S=\frac{1}{2}$ alternating FM-AFM dimerized spin chain realized in a distorted honeycomb lattice are Na3Cu2SbO6, Li3Cu2SbO6, and Na2Cu2TeO6. It should be noted that, despite the similarity in their crystal structures, spin chains in Li3Cu2SbO6 are fragmented due to a considerable site interchange between magnetic Cu2+ and non-magnetic Li1+ ions [Skakle1997, Nalbandyan2013, Koo2016, Vavilova2021]. We further note that a related compound Li2Cu2TeO6 has also been synthesized [Kumar2013] but its magnetic characterization still awaits studies. The crystal structure of these compounds can be viewed as a stacking of Na1+ layers alternating with Cu2MMO6 layers where $M=\mathrm{Sb},\mathrm{Te}$ [Smirnova2005, Xu2005]. A Cu2MMO6 layer is built upon edge-sharing elongated CuO6 octahedra forming a honeycomb lattice of Cu2+ ions with $M$ ions positioned at the center of hexagons. The honeycomb lattice is distorted due to a strong Jahn-Teller effect on oxygen octahedra of Cu2+ ions. Magnetic characteristics of these compounds are extensively studied by means of experiments as well as first-principles calculations. Experimental investigations include, for instance, inelastic neutron scattering (INS) [Miura2008, Gao2020, Shangguan2021], 23Na nuclear magnetic resonance (NMR) [Kuo2012], and thermodynamics measurements [Xu2005, Miura2006, Derakhshan2007, Schmitt2014, Patil2025]. The experimental studies reveal a finite energy gap with the non-magnetic singlet ground state in these systems. In particular, INS experiments unambiguously show that, at base temperatures, dimers are formed between third-nearest-neighbor rather than first-nearest-neighbor spins (see Fig. 1(c)). Surprisingly, fitting the energy dispersion of the triplet excitations yields minimal AFM $J_{3}$ coupling and FM $J_{2}$ coupling that are weaker than AFM $J_{1}$ even though the bond length of $J_{1}$ is longer than that of $J_{2}$.
@@
First-principles calculations employing density functional theory (DFT) [Xu2005, Derakhshan2007, Koo2008, Schmitt2014, Lin2022] and density-matrix renormalization group (DMRG) [Lin2022] provide a comprehensive microscopic description for such a peculiar behavior of these exchange couplings deduced from experiments. The DFT results suggest that the magnetism of these compounds is primarily governed by the half-filled Cu2+ $3d_{x^{2}-y^{2}}$ orbital, where the local $x$- and $y$-axis almost align along Cu2+-O2- bonds in the equatorial plane of the octahedron. This explains naturally the subordinate magnitude of $J_{3}$ coupling since the Cu2+-O2--Cu2+ path for $J_{3}$ incorporates a Cu2+-O2- bond (bond $B_{1}$ in Fig. 1(c)) that is elongated and perpendicular to the Cu2+ $3d_{x^{2}-y^{2}}$ orbital. That the magnitude of $J_{1}$ coupling is much greater than that of $J_{2}$ coupling owes to the direct overlap of Cu2+ $3d_{x^{2}-y^{2}}$ orbitals with O2- $2p$ orbitals along the Cu2+-O2--…-O2--Cu2+ path across the MMO6 octahedron. On the other hand, as displayed in Fig. 1(b), since the angle of the Cu2+-O2--Cu2+ path for $J_{2}$ is close to 90, a pair of O2- $2p$ orbitals is almost orthogonal resulting in a weaker $J_{2}$. In accordance with Goodenough-Kanamori-Anderson rules [Anderson1950, Goodenough1955, Goodenough1958, Moskvin1975], the sign of $J_{2}$ coupling is expected to be FM.
@@
Another compound in the family for exploring physics of the quasi-1D alternating dimerized spin chains has been reported in Ref. [Climent2011]. The authors of Ref. [Climent2011] have synthesized a Delafossite-derived compound Cu5SbO6 ($^{1+}_{3}Cu$$^{2+}_{2}$SbO6) that realizes a distorted honeycomb lattice of Cu2+ spins. The crystal structure of Cu5SbO6 is built from a monoclinic unit cell, as shown in Fig. 1(a), with the space group $C2/c$ (no. 15) and the lattice parameters $a$ = 8.923 Å, $b$ = 5.593 Å, $c$ = 11.845 Å, and $\beta$ = 103.585. The structure can also be viewed as a stacking of magnetic Cu2SbO6 layers alternating with non-magnetic Cu1+ layers. According to the measurements performed in Ref. [Climent2011], powder sample of Cu5SbO6 develops stacking faults which manifest as diffusive streaks along the $c^{*}$-axis in a selected-area electron diffraction pattern. The stacking faults can be effectively improved by elevating synthesis temperature. The temperature dependence of magnetic susceptibility and magnetic heat capacity indicates a formation of spin dimers as temperature is lowered below 120 K. An approximated energy gap of $\sim$190 K is obtained by fitting susceptibility data with an isolated dimer model. While it is plausible to assume a similar $J_{1}$-$J_{2}$-$J_{3}$ magnetic coupling scheme for Cu5SbO6, a DFT calculation [SaeFu2025] suggests a slightly different model. First of all, DFT calculation confirms the alternation of AFM $J_{1}$ and FM $J_{2}$ as in the case for related compounds. However, it is found that the AFM inter-plane coupling $J_{4}$ (see Fig. 1(d)) is significant, and $J_{3}$ is even smaller than in the related compounds. By inspecting the crystal structures of Cu5SbO6 and Na2Cu2TeO6 in between the honeycomb planes, as shown in Fig. 1(e), one can see that the two angles Cu2+-O2--Cu1+ of the super-exchange path Cu2+-O2--Cu1+-O2--Cu2+ in Cu5SbO6 amount to about 120 which facilitates the electron exchange more efficiently between the two Cu2+ ions compared to Na2Cu2TeO6 where the angles Cu2+-O2--Na1+ are close to 90.
@@ II.2 Magnetometry and neutron spectroscopy
Magnetization of Cu5SbO6 in an applied magnetic field oriented parallel ($\boldsymbol{H}_{ab}$) and perpendicular ($\boldsymbol{H}_{c^{*}}$) to the honeycomb plane was measured using a Quantum Design MPMS 3 vibrating-sample magnetometer. The temperature dependence was measured from $T=2$ to 400 K with $H=100$ Oe whereas the field dependence was measured from $H=0$ to $7\times 10^{4}$ Oe at $T=2$ and 120 K. For both field directions, a sample with mass 9.48 mg was aligned and attached to a quartz rod whose magnetization was measured separately under the same conditions.
@@ III.1 Crystal structure and stacking faults
Figure 3: (a) A projection of the crystal structure onto the $bc^{*}$ plane showing the ideal stacking sequence of Cu5SbO6. (b) Two equivalent representatives of in-plane translations of a honeycomb layer that contain one stacking fault depicted by the purple arrows. The grey arrows represent the ideal translation vector. ‘1’, ‘2’, and ‘3’ are sub-lattice labels of the honeycomb layer. The atom representations are the same as in Fig. 1(a). (c) A TEM image in the $ac$ plane of the sample synthesized at 1000 C. (d), (e) SAED patterns in the $(H,0,L)$ planes of samples synthesized at 1000 C and 1150 C, respectively. (f) A simulated SAED pattern computed with $p_{1}=p_{2}=1/2$. The crystal structure illustrations are generated by VESTA [Momma2011].
@@
Refinements of the crystal structure using the collected single-crystal and powder XRD data are based on the structural model reported in Ref [Climent2011]. The refinement with the single-crystal XRD data is performed with 799 unique reflections with $I_{\mathrm{obs}}>3\sigma$, where $I_{\mathrm{obs}}$ and $\sigma$ are the observed intensity and corresponding standard deviation, respectively, using jana2020 software package [Petricek2014]. The refinement result is presented in Fig. 2(a) as a plot of observed structure factor $F_{\mathrm{obs}}$ against calculated structure factor $F_{\mathrm{cal}}$ computed from the refined atomic parameters listed in Table 1. As can be seen in Fig. 2(a), the model provides the values of the calculated structure factor that are fairly consistent with the observed data. The cleavage of a typical crystal was examined with the $\theta$-$2\theta$ scan and was found to be $(00L)$ planes as shown in the inset of Fig. 2(a). Fig. 2(b) displays the powder XRD pattern obtained from ground crystals. Comparing the position of the observed Bragg peaks with that derived from the model (green bars) indicates that impurities in the sample, if present, are negligible. The diffraction pattern is fitted to the model using the Rietveld method [Rietveld1969] as implemented in fullprof software suite [Rodriguez-Carvajal1993]. The refinement yields the optimized lattice and atomic parameters listed in Table 1. The calculated diffraction pattern generated from the so-obtained parameters is illustrated in Fig. 2(b) as a black solid line which fits the observed data fairly well.
@@
While XRD is one of the most powerful techniques for investigating the periodic structure of compounds, it is, however, practically incapable of distinguishing ions that have similar atomic factors as in the case of Cu1+ and Cu2+ in Cu5SbO6. Namely, while coordinates of Cu(1), Cu(2) (between honeycomb planes), and Cu(3) (in honeycomb planes) in the unit cell can be accurately identified, their oxidation states must be determined by another method. As Cu1+ and Cu2+ are quite similar in size, it is possible to have site interchange between them as in the case of Li3Cu2SbO6 in which considerable amounts of Li1+ and Cu2+ ions exchange sites [Koo2016, Vavilova2021]. To confirm that all Cu(3) in Cu5SbO6 has 2+ oxidation state and, therefore, form full Cu2+ honeycomb layers, we turn to XAFS technique which is more effective for determining the oxidation state of a specific ion. The absorption coefficient, $\mu x$, of a material can be obtained from $I=I_{0}e^{-\mu x}\Leftrightarrow\mu x=-\ln(I/I_{0})$ where $I$ and $I_{0}$ are transmitted and incident intensities, respectively. Fig. 2(c) and 2(d) show the collected XAFS spectra of CuO, Cu2O and Cu5SbO6, respectively. Inspection on the results in Fig. 2(c) shows that Cu1+ ions in Cu2O (green) and Cu2+ ions in CuO (orange) have unequal edge energies. Estimated values of edge $E_{0}$ for Cu1+ and Cu2+ ions are determined from their absorption spectra as the energy $E$ at which the first derivative of $\mu x$ is maximized, which is done using the athena module in the demeter software suite [Ravel2005]. The first derivatives of $\mu x(E)$ for both Cu2O and CuO are displayed as dashed lines in Fig. 2(c) with the asterisk and diamond symbols marking the estimated $E_{0}$ of Cu1+ and Cu2+ ions, respectively. The difference in $E_{0}$ of the two ions is about 10 eV.
@@
To determine the oxidation states of Cu ions in Cu5SbO6, we performed the extended x-ray absorption fine structure (EXAFS) analysis. EXAFS analysis is performed in the oscillating region of the spectrum [Newville2014] which is quantified by $\varphi(k)\equiv\mu x-\mu_{0}x$ where $\mu_{0}x$ is the background signal (dashed line in Fig. 2(d)) and $k$ is related to $E$ through $k=\sqrt{2m_{\mathrm{e}}(E-E_{0})}/\hbar$. Here $m_{\mathrm{e}}$ is the electron mass. Contrary to XRD that depends upon average periodic structures, EXAFS relies on local environments around a particular ion. As the local environment of Cu sites within the honeycomb planes and that in between the planes are completely different, each site, therefore, produces a distinct oscillating pattern in $\varphi(k)$. To construct a model for the fitting, the crystal structure listed in Table 1 is used to generate all possible electron scattering paths around each Cu site, and $\varphi(k)$ of Cu5SbO6 is then a superposition of $\varphi_{i}(k)$’s from paths having significant amplitudes. To facilitate the fitting, the observed spectrum is Fourier-transformed to a real-space variable $R$, whose modulus and real part are illustrated in Figs. 2(e) and 2(f), respectively. The fitting is performed using the artemis module in the in demeter software suite [Ravel2005]. The fitting parameters include path distance corrections $\delta R$ along with thermal deviations $\sigma$, Cu amplitude reduction factor $S_{0}^{2}$, and edge corrections $\delta E_{0}$. The amplitude reduction factor $S_{0}^{2}$ for Cu atom is obtained separately from fitting $\varphi(R)$ collected with a Cu foil sample and is found to be 0.89(5). The initial values of $E_{0}$ for each Cu site are set to the value of Cu1+ ion, namely, the value marked by the asterisk in Fig. 2(c). The fitting result is shown as red curves in Figs. 2(e) and 2(f) with $R=0.262\%$. The optimized values of $\delta E_{0}$ for Cu(1) and Cu(2) are found to be the same (and, therefore, are constrained to the same value in the final refinement) which equals to 3.0(5) eV, while $\delta E_{0}$ of Cu(3) is 13.2(3) eV. Hence, $E_{0}$ of Cu(3) residing in the honeycomb planes is larger than $E_{0}$ of the other two sites in between the planes by about 10 eV, i.e., the oxidation state of Cu(3) is 2+ and that of Cu(1) and Cu(2) is 1+.

### v2

@@ Abstract
We analyze powder-averaged inelastic neutron scattering and magnetization data for the distorted honeycomb compound Cu5SbO6 using a first-order dimer expansion calculation and quantum Monte Carlo simulations. We show that, in contrast to the previously proposed honeycomb lattice model, Cu5SbO6 accommodates interacting dimerized spin chains with alternating ferromagnetic-antiferromagnetic couplings along the chain. Moreover, unlike the typical couplings observed in other Cu2+-based distorted honeycomb magnets, the spin chains in Cu5SbO6 primarily couple through an antiferromagnetic coupling that arises between the honeycomb layers, rather than the expected interchain coupling in the layers. This finding reveals a different magnetic coupling scheme for Cu5SbO6. In addition, utilizing x-ray spectroscopy and transmission electron microscopy, we also refine the crystal structure and stacking-fault model of the compound.
@@ I Introduction
Dimerized quantum magnets have constituted a fertile ground for exploring a plethora of exotic emergent quantum phenomena in magnetic insulators [1, 2]. The most renowned aspect of such magnets is arguably quantum criticality. One of the quantum phase transitions is achieved in the field-induced Bose-Einstein condensation of magnons realized in several compounds [3, 4, 5, 6, 7, 8, 9, 10]. Novel phenomena in the vicinity of a quantum critical point separating a dimerized disordered phase from a long-range ordered phase such as dimensional reduction [11] and longitudinal excitation modes [12] have been observed. When frustration between exchange couplings among orthogonally oriented dimers are present as in the Shastry-Sutherland magnet, several other intriguing effects have been reported including topologically protected chiral edge modes of triplet excitations [13], pressure-induced 4-spin plaquette singlet state [14, 15], field-induced spin-nematic phase [16], incomplete devil’s staircase [17] and crystallization of triplet bound states in magnetization plateaus [18]. Dimerized magnets are built upon spin dimers consisting of a pair of spins interacting with an intra-dimer antiferromagnetic (AFM) exchange coupling. The dimer, for the case of $S=\frac{1}{2}$, exhibits a non-magnetic singlet ground state and degenerate triplet excited states in zero magnetic field which are separated by a finite energy gap. Dimers that interact with each other with a ferromagnetic (FM) exchange coupling in a one-dimensional (1D) network, forming an alternating FM-AFM spin chain, form another distinct class of dimerized magnets. Magnetic materials that are characterized as quasi-1D FM-AFM dimerized spin chains include, as examples, copper nitrate [19], BaCu2V2O8 [20], CuNb2O6 [21], (CH3)2NH2CuCl3 [22], etc.
@@
Figure 1: (a) A monoclinic unit cell of Cu5SbO6 is displayed in which spin chains consisted of SbO6 octahedra (in silver) and edge-sharing Cu2O6 double plaquettes (in green) are aligned along the $a$-axis. The unit cell boundary is indicated by the dashed lines. (b) A chain segment with alternating $J_{1}$ and $J_{2}$ couplings is shown. The dashed purple and solid yellow lines represent intra-dimer $J_{1}$ and inter-dimer $J_{2}$ couplings, respectively. (c) The abab-plane projection of a single Cu2SbO6 layer is displayed. $B_{1}$ = 2.314 Å is the Cu-O bond perpendicular to the plaquette whereas $B_{2}$ = 1.994 Å and $B_{3}$ = 2.043 Å are the Cu-O bonds within the plaquette. The dot-dash turquoise lines represent inter-chain $J_{3}$ couplings. (d) A single layer projected on to the acac-plane is shown, where the dotted orange lines represent inter-chain $J_{4}$ couplings perpendicular to the honeycomb planes. The super-exchange path responsible for the $J_{4}$ coupling in Cu5SbO6 is displayed in (e) along with the corresponding path in Na2Cu2TeO6. The crystal structure illustrations are generated by VESTA [23].
@@
To our knowledge, the archetypes of the quasi-1D $S=\frac{1}{2}$ alternating FM-AFM dimerized spin chain realized in a distorted honeycomb lattice are Na3Cu2SbO6 [24, 25, 26, 27, 28], Li3Cu2SbO6 [29, 30], and Na2Cu2TeO6 [26, 28, 31, 32, 33]. It should be noted that, despite the similarity in their crystal structures, spin chains in Li3Cu2SbO6 are fragmented due to a considerable site interchange between magnetic Cu2+ and non-magnetic Li1+ ions [34, 35, 29, 30]. We further note that a related compound Li2Cu2TeO6 has also been synthesized [36] but its magnetic characterization still awaits studies. The crystal structure of these compounds can be viewed as a stacking of Na1+ layers alternating with Cu2MMO6 layers where $M=\mathrm{Sb},\mathrm{Te}$ [37, 38, 39, 40]. A Cu2MMO6 layer is built upon edge-sharing elongated CuO6 octahedra forming a honeycomb lattice of Cu2+ ions with $M$ ions positioned at the center of hexagons. The honeycomb lattice is distorted due to a strong Jahn-Teller effect on oxygen octahedra of Cu2+ ions. Magnetic characteristics of these archetypal compounds are extensively studied by means of experiments as well as first-principles calculations. Experimental investigations include, for instance, inelastic neutron scattering (INS) [25, 31, 32], 23Na nuclear magnetic resonance (NMR) [41, 27], and thermodynamics measurements [38, 24, 42, 28, 43]. The experimental studies reveal a finite energy gap with the non-magnetic singlet ground state in these systems. In particular, INS experiments unambiguously show that, at base temperatures, dimers are formed between third-nearest-neighbor rather than first-nearest-neighbor spins. Surprisingly, fitting the energy dispersion of the triplet excitations yields minimal AFM $J_{3}$ coupling and FM $J_{2}$ coupling that are weaker than AFM $J_{1}$ even though the bond length of $J_{1}$ is longer than that of $J_{2}$. Here, as indicated in Figs. 1(b) and 1(c), $J_{2}$ and $J_{3}$ denote the couplings between the first-nearest-neighbor spins whereas $J_{1}$ denotes the coupling between the third-nearest-neighbor spins.
@@
First-principles calculations employing density functional theory (DFT) [38, 42, 26, 28, 33] and density-matrix renormalization group (DMRG) [33] provide a comprehensive microscopic description for such a peculiar behavior of these exchange couplings deduced from experiments. The DFT results suggest that the magnetism of these compounds is primarily governed by the half-filled Cu2+ $3d_{x^{2}-y^{2}}$ orbital, where the local $x$- and $y$-axis almost align along Cu2+-O2- bonds in the equatorial plane of the octahedron. This explains naturally the subordinate magnitude of $J_{3}$ coupling since the Cu2+-O2--Cu2+ path for $J_{3}$ incorporates a Cu2+-O2- bond (bond $B_{1}$ in Fig. 1(c)) that is elongated and perpendicular to the Cu2+ $3d_{x^{2}-y^{2}}$ orbital. That the magnitude of $J_{1}$ coupling is much greater than that of $J_{2}$ coupling owes to the direct overlap of Cu2+ $3d_{x^{2}-y^{2}}$ orbitals with O2- $2p$ orbitals along the Cu2+-O2--…-O2--Cu2+ path across the MMO6 octahedron. On the other hand, as displayed in Fig. 1(b), since the angle of the Cu2+-O2--Cu2+ path for $J_{2}$ is close to 90, a pair of O2- $2p$ orbitals is almost orthogonal resulting in a weaker $J_{2}$. In accordance with Goodenough-Kanamori-Anderson rules [44, 45, 46, 47], the sign of $J_{2}$ coupling is expected to be FM.
@@
Another compound in the family for exploring physics of the quasi-1D alternating dimerized spin chains has been reported in Ref. [48]. The authors of Ref. [48] have synthesized a Delafossite-derived compound Cu5SbO6 ($^{1+}_{3}Cu$$^{2+}_{2}$SbO6) that realizes a distorted honeycomb lattice of Cu2+ spins. The crystal structure of Cu5SbO6 is built from a monoclinic unit cell, as shown in Fig. 1(a), with the space group $C2/c$ (no. 15) and the lattice parameters $a$ = 8.923 Å, $b$ = 5.593 Å, $c$ = 11.845 Å, and $\beta$ = 103.585. The structure can also be viewed as a stacking of magnetic Cu2SbO6 layers alternating with non-magnetic Cu1+ layers. According to the measurements performed in Ref. [48], powder sample of Cu5SbO6 develops stacking faults which manifest as diffusive streaks along the $c^{*}$-axis in a selected-area electron diffraction pattern. The stacking faults can be effectively improved by elevating synthesis temperature. The temperature dependence of magnetic susceptibility and magnetic heat capacity indicates a formation of spin dimers as temperature is lowered below 120 K. An approximated energy gap of $\sim$190 K is obtained by fitting susceptibility data with an isolated dimer model. While it is plausible to assume a similar $J_{1}$-$J_{2}$-$J_{3}$ magnetic coupling scheme for Cu5SbO6, a DFT calculation [49] suggests a slightly different model. First of all, DFT calculation confirms the alternation of AFM $J_{1}$ and FM $J_{2}$ as in the case for related compounds. However, it is found that the AFM inter-plane coupling $J_{4}$ (see Fig. 1(d)) is significant, and $J_{3}$ is even smaller than in the related compounds. By inspecting the crystal structures of Cu5SbO6 and Na2Cu2TeO6 in between the honeycomb planes, as shown in Fig. 1(e), one can see that the two angles Cu2+-O2--Cu1+ of the super-exchange path Cu2+-O2--Cu1+-O2--Cu2+ in Cu5SbO6 amount to about 120 which facilitates the electron exchange more efficiently between the two Cu2+ ions compared to Na2Cu2TeO6 where the angles Cu2+-O2--Na1+ are close to 90.
@@ II.2 Magnetometry and neutron spectroscopy
Magnetization of Cu5SbO6 in an applied magnetic field oriented parallel ($\boldsymbol{H}_{ab}$) and perpendicular ($\boldsymbol{H}_{c^{*}}$) to the honeycomb plane was measured using a Quantum Design MPMS 3 vibrating-sample magnetometer. The temperature dependence was measured from $T=2$ to 400 K with $H=100$ Oe whereas the field dependence was measured from $H=0$ to $7\times 10^{4}$ Oe at $T=2$ and 120 K. For both field directions, an aligned crystal sample (see inset of Fig. 2(a)) with mass 9.48 mg was attached to a quartz rod whose magnetization was measured separately under the same conditions.
@@ III.1 Crystal structure and stacking faults
Figure 3: (a) A projection of the crystal structure onto the $bc^{*}$ plane showing the ideal stacking sequence of Cu5SbO6. (b) Two equivalent representatives of in-plane translations of a honeycomb layer that contain one stacking fault depicted by the purple arrows. The grey arrows represent the ideal translation vector. ‘1’, ‘2’, and ‘3’ are sub-lattice labels of the honeycomb layer. The atom representations are the same as in Fig. 1(a). (c) A TEM image in the $ac$ plane of the sample synthesized at 1000 C. (d), (e) SAED patterns in the $(H,0,L)$ planes of samples synthesized at 1000 C and 1150 C, respectively. (f) A simulated SAED pattern computed with $p_{1}=p_{2}=1/2$. The crystal structure illustrations are generated by VESTA [23].
@@
Refinements of the crystal structure using the collected single-crystal and powder XRD data are based on the structural model reported in Ref [48]. The refinement with the single-crystal XRD data is performed with 799 unique reflections with $I_{\mathrm{obs}}>3\sigma$, where $I_{\mathrm{obs}}$ and $\sigma$ are the observed intensity and corresponding standard deviation, respectively, using jana2020 software package [53]. The refinement result is presented in Fig. 2(a) as a plot of observed structure factor $F_{\mathrm{obs}}$ against calculated structure factor $F_{\mathrm{cal}}$ computed from the refined atomic parameters listed in Table 1. As can be seen in Fig. 2(a), the model provides the values of the calculated structure factor that are fairly consistent with the observed data. The cleavage of a typical crystal was examined with the $\theta$-$2\theta$ scan and was found to be $(00L)$ planes as shown in the inset of Fig. 2(a). Fig. 2(b) displays the powder XRD pattern obtained from ground crystals. Comparing the position of the observed Bragg peaks with that derived from the model (green bars) indicates that impurities in the sample, if present, are negligible. The diffraction pattern is fitted to the model using the Rietveld method [54] as implemented in fullprof software suite [55]. The refinement yields the optimized lattice and atomic parameters listed in Table 1. The calculated diffraction pattern generated from the so-obtained parameters is illustrated in Fig. 2(b) as a black solid line which fits the observed data fairly well.
@@
While XRD is one of the most powerful techniques for investigating the periodic structure of compounds, it is, however, practically incapable of distinguishing ions that have similar atomic factors as in the case of Cu1+ and Cu2+ in Cu5SbO6. Namely, while coordinates of Cu(1), Cu(2) (between honeycomb planes), and Cu(3) (in honeycomb planes) in the unit cell can be accurately identified, their oxidation states must be determined by another method. As Cu1+ and Cu2+ are quite similar in size, it is possible to have site interchange between them as in the case of Li3Cu2SbO6 in which considerable amounts of Li1+ and Cu2+ ions exchange sites [29, 30]. To confirm that all Cu(3) in Cu5SbO6 has 2+ oxidation state and, therefore, form full Cu2+ honeycomb layers, we turn to XAFS technique which is more effective for determining the oxidation state of a specific ion. The absorption coefficient, $\mu x$, of a material can be obtained from $I=I_{0}e^{-\mu x}\Leftrightarrow\mu x=-\ln(I/I_{0})$ where $I$ and $I_{0}$ are transmitted and incident intensities, respectively. Fig. 2(c) and 2(d) show the collected XAFS spectra of CuO, Cu2O and Cu5SbO6, respectively. Inspection on the results in Fig. 2(c) shows that Cu1+ ions in Cu2O (green) and Cu2+ ions in CuO (orange) have unequal edge energies. Estimated values of edge $E_{0}$ for Cu1+ and Cu2+ ions are determined from their absorption spectra as the energy $E$ at which the first derivative of $\mu x$ is maximized, which is done using the athena module in the demeter software suite [56]. The first derivatives of $\mu x(E)$ for both Cu2O and CuO are displayed as dashed lines in Fig. 2(c) with the asterisk and diamond symbols marking the estimated $E_{0}$ of Cu1+ and Cu2+ ions, respectively. The difference in $E_{0}$ of the two ions is about 10 eV.
@@
To determine the oxidation states of Cu ions in Cu5SbO6, we performed the extended x-ray absorption fine structure (EXAFS) analysis. EXAFS analysis is performed in the oscillating region of the spectrum [57] which is quantified by $\varphi(k)\equiv\mu x-\mu_{0}x$ where $\mu_{0}x$ is the background signal (dashed line in Fig. 2(d)) and $k$ is related to $E$ through $k=\sqrt{2m_{\mathrm{e}}(E-E_{0})}/\hbar$. Here $m_{\mathrm{e}}$ is the electron mass. Contrary to XRD that depends upon average periodic structures, EXAFS relies on local environments around a particular ion. As the local environment of Cu sites within the honeycomb planes and that in between the planes are completely different, each site, therefore, produces a distinct oscillating pattern in $\varphi(k)$. To construct a model for the fitting, the crystal structure listed in Table 1 is used to generate all possible electron scattering paths around each Cu site, and $\varphi(k)$ of Cu5SbO6 is then a superposition of $\varphi_{i}(k)$’s from paths having significant amplitudes. To facilitate the fitting, the observed spectrum is Fourier-transformed to a real-space variable $R$, whose modulus and real part are illustrated in Figs. 2(e) and 2(f), respectively. The fitting is performed using the artemis module in the in demeter software suite [56]. The fitting parameters include path distance corrections $\delta R$ along with thermal deviations $\sigma$, Cu amplitude reduction factor $S_{0}^{2}$, and edge corrections $\delta E_{0}$. The amplitude reduction factor $S_{0}^{2}$ for Cu atom is obtained separately from fitting $\varphi(R)$ collected with a Cu foil sample and is found to be 0.89(5). The initial values of $E_{0}$ for each Cu site are set to the value of Cu1+ ion, namely, the value marked by the asterisk in Fig. 2(c). The fitting result is shown as red curves in Figs. 2(e) and 2(f) with $R=0.262\%$. The optimized values of $\delta E_{0}$ for Cu(1) and Cu(2) are found to be the same (and, therefore, are constrained to the same value in the final refinement) which equals to 3.0(5) eV, while $\delta E_{0}$ of Cu(3) is 13.2(3) eV. Hence, $E_{0}$ of Cu(3) residing in the honeycomb planes is larger than $E_{0}$ of the other two sites in between the planes by about 10 eV, i.e., the oxidation state of Cu(3) is 2+ and that of Cu(1) and Cu(2) is 1+.